OpenStudy (anonymous):

Find f(5) for this sequence f(1)=2 and f(2)=5 f(n)=f(1)+f(2)+f(n-1) for n>2 btw it doesn't let me type back

OpenStudy (anonymous):

what doesn't let you type back?

OpenStudy (anonymous):

\[f(1)=2,f(2)=5,f(3)=f(1)+f(2)=2+5=7\]

OpenStudy (anonymous):

then \[f(4)=f(1)+f(2)+f(3)=2+5+7=14\]

OpenStudy (anonymous):

one more and you are done